OpenStudy (anonymous):

3x-27= 52 5x+y=26 Solve using elimination for system of equations

OpenStudy (anonymous):

Is that 3x-27Y=52?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve for Y

OpenStudy (anonymous):

@Study23 yes, it is 27Y

OpenStudy (anonymous):

Okay, so eliminate one variable. I would suggest eliminating the x first. Do you know how to do that?

OpenStudy (anonymous):

3x-27= 52 5x+y=26 multiply first equation to 5 and second equation to 3

OpenStudy (anonymous):

so would 14 be the answer?

OpenStudy (anonymous):

or is it -14

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

sorry, thats not an option

OpenStudy (anonymous):

its 14 -14 -18 1 those are the options

OpenStudy (anonymous):

One moment...

OpenStudy (anonymous):

@selenamalter What is the question asking you to find?

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

@Study23 Y

OpenStudy (anonymous):

the value of y to satisfy the equation

OpenStudy (anonymous):

I have an answer... Just double checking

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

before double checking, whats your answer?

OpenStudy (anonymous):

i wanna see it if matcheds mine

OpenStudy (anonymous):

@Study23

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Still here

OpenStudy (anonymous):